Write the standard form of the equation of the line through the given point with the given slope.

Through: (1,2), slope 7

Found 2 solutions by Alan3354, ultrakimidis1:
Answer by Alan3354(69443)

Through: (1,2), slope 7
---------
Use y-y1 = m*(x-x1) where m = slope and (x1,y1) is the point.
----
convert to any form you like.

Answer by ultrakimidis1(8)

Standard form of writing equation ax+by+c=0
Given point (1,2) and gradient m=7
Using slope methode:
y=mx+c......(1)
Put x=1 and y =2 and m=7 into eqn (1)
2=7(1)+c
2=7+c
c=2-7
c=-5
Put c=-5 and m=7 into eqn (1)
We have;
y=7x-5
Into its standard form, We have...
y-7x+5=0